# IsoSEL: Isometric Structural Entropy Learning for Deep Graph Clustering in Hyperbolic Space

### v1

@@ VII-A Neural Transformation in Lorentz Group
Algorithm 1 Training IsoSEL
@@
0:  A weighted graph $G=(\mathcal{V},\mathcal{E},\mathbf{X})$ and the height of deep partitioning tree $H$.
@@
0:  The partitioning tree $\mathcal{T}$, and tree node embeddings $\{\mathbf{Z}^{h}\}_{h=1,...,H}$ at all the levels.
1:  Generate the augmented graph using Eq. 25;
2:  while not converged do
3:     for both original and augmented graph do
@@
4:        Obtain leaf node embeddings $\mathbf{Z}^{H}=\operatorname{LConv}(\mathbf{X},\mathbf{A})$;
@@
5:        Compute the level-wise assignment and parent node embeddings recursively with Eqs. 31, 33 and 35;
6:     end for
7:     Forward leaf node embeddings into Lorentz projection head with Eq. 37;
8:     Compute the overall objective $\mathcal{L}_{\operatorname{deeptree}}+\alpha\mathcal{L}_{\operatorname{contrast}}$;
@@
DEC 23.54$\pm$1.13 15.13$\pm$0.72 28.34$\pm$0.42 28.12$\pm$0.24 37.35$\pm$0.84 18.29$\pm$0.64 38.56$\pm$0.24 34.76$\pm$0.35 22.44$\pm$0.14 19.55$\pm$0.13
VGAE 43.40$\pm$1.62 37.50$\pm$2.13 32.70$\pm$0.30 33.10$\pm$0.55 66.01$\pm$0.84 56.24$\pm$0.22 37.62$\pm$0.24 22.16$\pm$0.16 30.61$\pm$1.71 30.15$\pm$1.23
ARGA 48.10$\pm$0.45 44.10$\pm$0.28 35.10$\pm$0.58 34.60$\pm$0.48 58.36$\pm$1.02 44.18$\pm$0.85 37.21$\pm$0.58 26.28$\pm$1.38 24.80$\pm$0.17 24.35$\pm$0.17
MVGRL 62.91$\pm$0.42 56.96$\pm$0.74 46.96$\pm$0.25 44.97$\pm$0.57 36.89$\pm$1.31 18.77$\pm$1.54 10.12$\pm$2.21     5.53$\pm$1.78 31.59$\pm$1.45 29.42$\pm$1.06
Sublime 54.20$\pm$0.28 50.32$\pm$0.32 44.10$\pm$0.27 43.91$\pm$0.35     6.37$\pm$0.54     5.36$\pm$0.24 39.16$\pm$1.82 24.15$\pm$0.63 32.01$\pm$1.24 30.36$\pm$0.85
ProGCL 41.02$\pm$1.64 30.71$\pm$1.38 39.59$\pm$1.58 36.16$\pm$2.21 39.56$\pm$0.25 34.18$\pm$0.58 35.50$\pm$2.06 26.08$\pm$1.91 30.24$\pm$1.21 28.98$\pm$1.06
GRACE 57.30$\pm$0.86 52.70$\pm$1.20 39.90$\pm$2.26 37.70$\pm$1.35 53.46$\pm$1.32 42.74$\pm$1.57 47.90$\pm$1.65 36.40$\pm$1.56 30.08$\pm$0.24 27.60$\pm$1.24
MCGC 44.90$\pm$1.56 37.80$\pm$1.24 40.14$\pm$1.44 38.00$\pm$0.85 61.54$\pm$0.29 52.10$\pm$0.27 53.17$\pm$1.29 39.02$\pm$0.53 33.39$\pm$0.02 29.25$\pm$0.01
DCRN 48.86$\pm$0.85 43.79$\pm$0.48 45.86$\pm$0.35 47.64$\pm$0.30 73.70$\pm$1.24 63.69$\pm$0.84 OOM OOM 32.20$\pm$0.08 31.41$\pm$0.12
FT-VGAE 61.03$\pm$0.52 58.22$\pm$1.27 44.50$\pm$0.13 46.71$\pm$0.75 69.76$\pm$1.06 59.30$\pm$0.81 51.36$\pm$0.92 40.07$\pm$2.13 33.05$\pm$0.24 35.50$\pm$0.12
gCooL 58.33$\pm$0.24 56.87$\pm$1.03 47.29$\pm$0.10 46.78$\pm$1.51 63.21$\pm$0.09 52.40$\pm$0.11 47.42$\pm$1.76 27.71$\pm$2.28 30.82$\pm$0.15 31.48$\pm$0.34
S3GC 58.90$\pm$1.81 54.40$\pm$2.52 44.12$\pm$0.90 44.80$\pm$0.65 59.78$\pm$0.45 56.13$\pm$0.58 54.80$\pm$1.22 29.93$\pm$0.22 33.32$\pm$0.34 34.50$\pm$0.57
Congregate 63.16$\pm$0.71 59.27$\pm$1.23 50.92$\pm$1.58 47.59$\pm$1.60 70.99$\pm$0.67 60.55$\pm$1.36 46.03$\pm$0.47 38.57$\pm$1.05 35.03$\pm$0.23 33.42$\pm$0.47
DinkNet 62.28$\pm$0.24 61.61$\pm$0.90 45.87$\pm$0.24 46.96$\pm$0.30 74.36$\pm$1.24 68.40$\pm$1.37 39.54$\pm$1.52 33.87$\pm$0.12 36.39$\pm$0.34 35.77$\pm$0.13
GC-Flow 62.15$\pm$1.35 63.14$\pm$0.80 40.50$\pm$1.32 42.62$\pm$1.44 36.45$\pm$1.21 37.24$\pm$1.26 41.10$\pm$1.06 35.60$\pm$1.82 36.82$\pm$0.28 37.31$\pm$0.04
GCSEE 53.02$\pm$0.41 51.22$\pm$0.88 35.66$\pm$0.42 36.29$\pm$0.37 64.15$\pm$0.68 56.76$\pm$1.24 39.03$\pm$1.58 35.24$\pm$1.56 26.88$\pm$0.21 26.15$\pm$0.24
FastDGC 53.03$\pm$0.03 51.25$\pm$0.05 45.85$\pm$0.08 44.12$\pm$0.05 69.76$\pm$0.03 65.71$\pm$0.04 45.26$\pm$0.52 40.23$\pm$0.18 25.85$\pm$0.86 25.08$\pm$0.73
MAGI 59.72$\pm$1.22 57.31$\pm$1.31 45.28$\pm$0.24 46.86$\pm$0.47 71.62$\pm$0.11 61.55$\pm$0.24 54.22$\pm$1.88 46.27$\pm$1.64 33.28$\pm$0.04 31.29$\pm$0.08
RGC 57.60$\pm$1.36 50.46$\pm$1.72 45.70$\pm$0.29 45.47$\pm$0.43 47.65$\pm$0.91 42.65$\pm$1.53 46.24$\pm$1.05 36.12$\pm$0.30 28.01$\pm$1.85 25.56$\pm$1.91
DGCluster 61.47$\pm$0.34 44.75$\pm$0.28 43.75$\pm$0.05 31.55$\pm$0.09 77.12$\pm$0.25 67.59$\pm$0.35 50.10$\pm$1.87 38.92$\pm$1.27 33.73$\pm$0.68 21.84$\pm$0.71
IsoSEL 64.56$\pm$0.72 64.02$\pm$0.70 53.18$\pm$1.11 50.45$\pm$1.41 74.54$\pm$0.44 69.24$\pm$0.52 55.93$\pm$0.53 43.67$\pm$0.35 36.91$\pm$0.33 38.52$\pm$0.23
Avg. Gain 3.51 8.58 7.70 5.45 5.82 8.91 1.36 1.60 1.89 7.77
VII-B Hyperbolic Reweighing for Imbalance Clusters
In deep clustering on imbalanced graphs, minority clusters of fewer nodes are typically underrepresented due to the limited number of samples, which often results in a biased boundary among imbalanced clusters. To address this issue, in Lorentz tree contrastive learning, we propose hyperbolic reweighed sampling to mitigate the bias in clustering boundary. The core idea of reweighed sampling is to refine the hyperbolic representations of the majority clusters in the pushing-pulling process, as the structural entropy has a promising ability to identify minority clusters, as shown in Theorem IV.3.
To be specific, we first feed the original and augmented graphs into LSEnet to generate two hyperbolic views with their representations denoted as $\boldsymbol{z}$ and $\hat{\boldsymbol{z}}$, respectively. Second, we consider the contrast among the leaf nodes as the unity of the partitioning tree is ensured by the Lorentz projection head (Theorem VII.2). For each leaf node $\boldsymbol{z}^{H}_{i}$ in the original view, the positive sample is given as the corresponding $\hat{\boldsymbol{z}}^{H}_{i}$ in the augmented view. As for the negative samples, we draw $m$ leaf nodes of the same ancestor in the partitioning tree using a weighting function defined as follows,
$w_{ij}=\exp\left(-d_{\mathbb{L}}(L(\boldsymbol{z}^{H}_{i}),L(\hat{\boldsymbol{z}}^{H}_{j}))\right).\vspace{-0.05in}$ (39)
Accordingly, the Lorentz tree contrastive loss is formulated as
where $\tau$ is the temperature to control the strictness for negative samples, and $L$ denotes the Lorentz projection head. Note that, the reweighed sampling alleviates the over aggregation of majority cluster representations while preserving the closeness of similar nodes in hyperbolic space.
VII-C Overall Objective for Deep Graph Clustering
Finally, the overall objective of IsoSEL is formulated as
$\mathcal{J}=\mathcal{L}_{\operatorname{deeptree}}+\alpha\mathcal{L}_{\operatorname{contrast}},\vspace{-0.05in}$ (41)
where $\mathcal{L}_{\operatorname{deeptree}}$ is defined by Eq. (17), and $\alpha$ denotes the weighting coefficient. The training procedure is described in Algorithm 1, where Lorentz tree contrastive learning refines the deep partitioning tree created by differentiable structural information in hyperbolic space to generate promising node clusters. Specifically, we first learn the partitioning tree representation in a bottom-up fashion, as outlined in Lines $3$-$6$ of Algorithm 1. Then, we recover the tree through a top-down process in Lines $10$-$17$, ensuring that empty leaf nodes are avoided and single chains are fixed in the relaxation. Further details of the recovery procedure are provided in Algorithm 2 given in Appendix III.
Computational Complexity. In IsoSEL, the computational complexity of learning the deep partitioning tree is $\mathcal{O}(N)$ thanks to the node-wise formulation of differential structural entropy in Eq. (17). The Lorentz tree contrastive learning costs $\mathcal{O}(mN)$, where $N$ and $m$ denotes the number of nodes and negative samples, respectively. Accordingly, the overall complexity is yielded as $\mathcal{O}(N+mN)$, which is significantly less than that of our previous work of $\mathcal{O}(N|\mathcal{E}|)$ [20]. Note that, though the number of clusters $K$ is no longer required, IsoSEL presents competitive complexity to typical deep clustering models of $\mathcal{O}(KN)$.
In summary, IsoSEL takes the advantages of both hyperbolic geometry and differential structural entropy to effectively discover the cluster structure without $K$, while alleviating the clustering bias among imbalanced clusters.
@@ VIII Experiments
In the experiment, we compare with $20$ strong baselines on $5$ commonly used datasets, aiming to answer the research questions as follows: RQ1 (Main Results): How does the proposed IsoSEL perform in terms of both effectiveness and efficiency? RQ2: How does IsoSEL outperform the classic structural entropy [16] and our previous model [20]? RQ3 (Ablation Study): How does each proposed module contribute to the success of IsoSEL, including the choice of hyperbolic space, augmentation strategy and tree contrastive learning? RQ4 (Hyperparameter Sensitivity): How do the hyperparameters affect the performance? RQ5: How do the imbalance clusters impact the clustering results? Besides, we examine the embedding expressiveness of IsoSEL, and include a visualization of the clusters learnt from IsoSEL. Additional results and further details are provided in Appendix VII. Codes and data are publicly available at https://github.com/RiemannGraph/DSE_clustering.

### v2

@@ VI-A Neural Lorentz Boost
VI-B Tree Contrastive Loss
In this part, we introduce the loss function for contrasting partitioning trees in the Lorentz model of hyperbolic space. For each node, the corresponding node in the counterpart view is treated as positive sample; otherwise, negative samples. Accordingly, Tree Contrastive Loss (TCL) is formulated as
$\mathcal{L}_{\text{TCL}}(\mathbf{Z},\mathbf{Z^{\prime}})=\frac{1}{N}\sum_{i=1}^{N}\frac{\exp(d_{\mathbb{L}}(L(\boldsymbol{z_{i}}),L(\boldsymbol{z^{\prime}_{i}})))}{\sum_{j=1}^{N}w_{ij}\exp(d_{\mathbb{L}}(L(\boldsymbol{z_{i}}),L(\boldsymbol{z^{\prime}_{j}})))},\vskip-5.05942pt$ (36)
where the reweighting factor is defined with the level-wise assignment, $w_{ij}=\sum_{k=1}^{N_{H-1}}S^{H}_{ik}S^{H}_{jk}$. That is, TCL considers the (hierarchical) tree branches in the partitioning trees, essentially differing itself from graph contrastive learning.
Differences to Graph Contrastive Learning (GCL). GCL refines representations through exploring the similarity among all nodes in the graph. On the contrary, we contrast the partitioning trees of the graph—nodes and their parent hierarchies—under consistency insurance, so that we propose the Neural Lorentz Boost with linearity and isometry, maintaining the partitioning tree as a whole, and Tree Contrastive Loss, considering the partitioning structures. Also, GCL is typically conducted in Euclidean space, while we operate in hyperbolic space given its alignment to tree structures.
Remaining Issues. Although tree contrastive learning refines node representations in hyperbolic space, it presents the following two limitations: On the one hand, view generation in this case is non-trivial. A straightforward way is to perform graph augmentation followed by LSEnet to generate a new hyperbolic view, but it may introduce noise that undermines model expressiveness as evidenced in Ablation Study. On the other hand, computing the loss function of Eq. (36) yields a quadratic complexity with respect to $|\mathcal{V}|$, the number of nodes, preventing the model scaling up to large graphs.
VII Augmented Structural Information Learning
In this section, we first elaborate on a theoretical insight that the structural entropy bounds contrastive learning, allowing for efficient and implicit tree contrastive learning. Integrating the theoretical establishment and neural net designs, we finally propose a novel deep graph clustering approach, Augmented Structural Information Learning (ASIL).
VII-A Theoretical Insight: Connection between Structural Entropy and Contrastive Learning
We uncover an inherent connection between structural entropy and contrastive learning, while simultaneously addressing the two aforementioned limitations.
Specifically, we demonstrate that the objective of structural entropy aligns closely with that of contrastive learning in terms of graph conductance, and this connection is formally established by introducing a virtual graph constructed from node representations. First, prior to the mathematical formulation, we provide an intuitive discussion of their relationship: in hyperbolic space, the tree contrastive learning pulls together node representations belonging to the same partition (i.e., tree branch), while pushing apart those from different partitions; the structural entropy operates directly on the graph structure—namely, the adjacency matrix—and generates an optimal partitioning tree primarily based on the node connectivity. Thus, both structural entropy and contrastive learning explore the node similarity, but from structural and representational perspectives, respectively. Second, we introduce a virtual graph constructed from node representations to bridge these two viewpoints. In particular, we perform the Lorentz Boost Construction defined as follows,
$\tilde{\mathbf{A}}_{ij}=\exp\left(-d_{\mathbb{L}}\left(L(\boldsymbol{z}_{i}),L(\boldsymbol{z}_{j})\right)/t\right),\vskip-5.78172pt$ (37)
where $d_{\mathbb{L}}$ denotes the induced distance of Minkowski metric in the hyperbolic space, and $t$ is the temperature hyperparameter. The transformation $L$ corresponds to the Lorentz boost within the tree contrastive learning, and $\boldsymbol{z}$ refers to the optimized node representation. In practice, we sparsify the virtual adjacency matrix by retaining $k$-nearest neighbors regarding $d_{\mathbb{L}}$. An interesting finding is formally stated in the theorem below.
{mymath}
Theorem VII.1 (Connection between Structural Entropy and Tree Contrastive Learning).
[…]
Proof.
Refer to Appendix II.I, Theorem VII.1. ∎
The theorem above claims that structural entropy indeed bounds the tree contrastive learning defined in Eq. (36). The key point is Eq. (36) has been proven to align with minimizing the graph conductance, which is bounded by structural entropy as formalized in Theorem III.5.
Our theoretical insight offers a promising alternative for tree contrastive learning, addressing its limitation in view generation and computational complexity. Concretely, it is preferable to perform structural information minimization over the virtual graph, which implicitly contrasts the partitioning trees in the hyperbolic space. The effect of contrasting two hyperbolic views is captured in the virtual graph of Lorentz boost construction in a parameterized fashion, and thus view generation is no longer required. As for computational complexity, we achieve the linear complexity regarding $|\mathcal{V}|$ with the reformulation in Eq. (17), compared to the quadratic complexity of contrastive learning in Eq. (36).
VII-B Augmented Structural Entropy: Integrating Partitioning Tree Construction and Contrastive Learning
This part presents our deep graph clustering approach —Augmented Structural Information Learning, referred to as ASIL. Based on the established theory of differential structural information and findings on its connection to tree contrastive learning, ASIL seamlessly integrates partitioning tree construction and (implicit) contrastive learning in the hyperbolic space, allowing for effective and efficient node clustering without the knowledge of the number of clusters.
Specifically, ASIL first fuses the structural and representational views, and then feed them into LSEnet (Sec. V) to learn the hyperbolic partitioning tree for clustering. The overall architecture is sketched in Fig. 1. The key design is a simple yet expressive clustering objective—Augmented Structural Entropy, which is formalized below
$\min\nolimits_{\mathbf{\Theta}}\ \mathcal{H}^{\mathcal{T}_{\text{net}}}(G^{\gamma};\mathbf{Z};\mathbf{\Theta}),$ (38)
where $G^{\gamma}$ is a fusion of the original graph and its representational view, the parameterized virtual graph (Sec. VII-A) that is constructed from node representations derived by Lorentz convolution in practice. Thus, the adjacency matrix is given with a balancing factor $\gamma$ as follows,
$\mathbf{A}^{\gamma}=(1-\gamma)\mathbf{A}+\gamma\tilde{\mathbf{A}},\quad\gamma>0.$ (39)
Note that the objective of Eq. (38) simultaneously learns an enhanced partitioning tree with a virtual graph, in which optimizing the virtual graph refines node representation via implicitly Lorentz tree contrastive learning (Sec. VI).
Interestingly, this enhanced partitioning tree enjoys superior clustering capability to the deep tree of $\min\nolimits_{\mathbf{\Theta}}\ \mathcal{H}^{\mathcal{T}_{\text{net}}}(G;\mathbf{Z};\mathbf{\Theta})$, which is proved in light of graph conductance.
{mymath}
Theorem VII.2 (Graph Conductance Improvement with Graph Fusion).
Given a graph $G=(\mathcal{V},\mathcal{E})$ and a virtual graph $\tilde{G}$ as in Theorem VII.1, the fused graph $G^{\gamma}$ is derived as in Eq. (39). For any nontrivial subset $\mathcal{S}\subset\mathcal{V}$ with $0<\mathrm{vol}_{A}(\mathcal{S})<\mathrm{vol}_{A}(V)$, denote the conductance of $\mathcal{S}$ under a graph $G$ as
$\phi_{G}(\mathcal{S})=\frac{\mathrm{cut}_{G}(\mathcal{S},\mathcal{V}\setminus\mathcal{S})}{\min\{\mathrm{vol}_{G}(\mathcal{S}),\,\mathrm{vol}_{G}(\mathcal{V}\setminus\mathcal{S})\}},$
where $\mathrm{cut}_{G}(\mathcal{S},\mathcal{P})=\sum_{i\in\mathcal{S},j\in\mathcal{P}}w_{ij}$ and $\mathrm{vol}_{G}(\mathcal{S})=\sum_{i\in\mathcal{S}}\sum_{j}w_{ij}$. Suppose that for every $\mathcal{S}\subset\mathcal{V}$,
$\phi_{\tilde{G}}(\mathcal{S})\leq\phi_{G}(\mathcal{S}).$ (40)
Then, the conductance of the fused graph satisfies
$\phi_{G^{\alpha}}(\mathcal{S})\leq\phi_{G}(\mathcal{S}),\quad\forall\mathcal{S}\subset V.$
Moreover, if there exists a subset $\mathcal{S}^{\star}$ such that $\phi_{\tilde{G}}(\mathcal{S}^{\star})<\phi_{G}(\mathcal{S}^{\star})$, then $\phi_{G^{\alpha}}(\mathcal{S}^{\star})<\phi_{G}(\mathcal{S}^{\star})$ for all $\alpha\in(0,1]$.
Proof.
Refer to Appendix II.I, Theorem VII.2. ∎
Algorithm 1 Training ASIL
@@ VII-B Augmented Structural Entropy: Integrating Partitioning Tree Construction and Contrastive Learning
0: A weighted graph $G=(\mathcal{V},\mathcal{E},\mathbf{X})$ and the height of deep partitioning tree $H$, fusion parameter $\gamma\in(0,1]$, nearset neighbors number $k$.
@@
0: The partitioning tree $\mathcal{T}$, and tree node embeddings $\{\mathbf{Z}_{\gamma}^{h}\}_{h=1,...,H}$ at all the levels.
1:while not converged do
2:  Obtain initial embeddings $\mathbf{Z}^{H}=\operatorname{LConv}(\mathbf{X},\mathbf{A})$;
3:  Infer virtual graph $\tilde{\mathbf{A}}_{ij}=\exp(-d_{\mathbb{L}}(L(\boldsymbol{z}_{i}),L(\boldsymbol{z}_{j}))/\tau)$ with Lorentz boost $L$ in Eq. 32;
4:  Sparsify the virtual graph with $k$-nearest neighbors;
5:  Conduct graph fusion by $\mathbf{A}^{\gamma}=\mathbf{A}+\gamma\tilde{\mathbf{A}}$;
@@
6:  Obtain leaf node embeddings $\mathbf{Z}_{\gamma}^{H}=\operatorname{LConv}(\mathbf{X},\mathbf{A}^{\gamma})$;
@@
7:  Compute the level-wise assignment and parent node embeddings $\{\mathbf{Z}_{\gamma}^{h}\}_{h=1,...,H-1}$ recursively with Eqs. 27, 29 and 31;
8:  Compute the overall objective $\mathcal{H}^{\mathcal{T}_{\text{net}}}(G^{\gamma};\mathbf{Z};\mathbf{\Theta})$;
@@
VII-C Algorithm: ASIL for Graph Clustering
The overall procedure of ASIL is summarized in Algorithm 3. The first phase (Line 1-10) is to learn hyperbolic partitioning tree representation in a bottom-up manner. Specifically, in Line 2-4, a parameterized virtual graph first is constructed from node representations with a neural Lorentz boost, implicitly capturing tree contrastive learning. In Line 2-4, we forward a fused graph into LSEnet, recursively inducing parent node embeddings. In Line 8-9, the partitioning tree is optimized by a novel objective of augmented structural entropy. The second phase (Line 11-18) is to recover the partitioning tree structure through a top-down decoding process, which prunes empty leaves and resolves redundant chains in accordance with Theorem IV.2. Clustering structures are naturally uncovered in the partitioning tree, whose procedure are provided in Algorithm 2 given in Appendix III.
Computational Complexity. First, we analyze the complexity of computing differentiable structural information. Thanks to the node-wise formulation in Eq. (17), this computation no longer requires enumerating all edges at each level. Instead, it operates solely on the coarsened graphs $\{G^{h}\}_{h=1}^{H}$, whose node counts $\{N_{h}\}$ satisfy $N_{h}\ll N_{h-1}$ and $N_{H}=N$. Second, given that the induced graph $G^{h-1}$ is sparse, $w^{h}_{kl}$ is computed via sparse matrix multiplication $\mathbf{W}^{h}=(\mathbf{S}^{h})^{\top}\mathbf{A}\mathbf{S}^{h}$, yielding the computational cost of $\mathcal{O}(|\mathcal{E}|)$ at leaf level, where $\mathcal{E}$ is the edge set of graph $G$. Third, we leverage $k$-nearest neighbors to construct the virtual adjacency matrix, which costs $\mathcal{O}(k|\mathcal{V}|)$, where $\mathcal{V}$ is the node set. Finally, the objective of augmented structural entropy regarding $\mathbf{A}^{\gamma}=(1-\gamma)\mathbf{A}+\gamma\tilde{\mathbf{A}}$ costs $\mathcal{O}(|\mathcal{E}|+k|\mathcal{V}|)$ in total. That is, the proposed ASIL discovers node clusters with a more preferable linear complexity in graph size, unknowing the number of clusters, while previous methods (e.g., MVGRL [21]) typically undergo a clustering procedure costing $\mathcal{O}(|\mathcal{N}|^{2})$.
In summary, we present a hyperbolic deep model ASIL built on the established theory of differential structural entropy that allows for effective and efficient debiased clustering in linear complexity, without requiring the number of clusters.
@@ VII-C Algorithm: ASIL for Graph Clustering
[…]
Avg. Gain 11.55 15.16 12.42 10.25 18.20 20.52 12.84 12.82 6.95 10.02
@@ VIII Experiments
In the experiment, we compare with $20$ strong baselines on $5$ commonly used datasets, aiming to answer the research questions as follows: RQ1 (Main Results): How does the proposed ASIL perform in terms of both effectiveness and efficiency, especially compared to vallina LSEnet? RQ2: How does ASIL outperform the classic structural entropy [27] and our previous model [54]? RQ3 (Ablation Study): How does each proposed module contribute to the success of ASIL, including the choice of hyperbolic space, Neural Lorentz Boost Construction and Implicit Tree Contrastive Learning? RQ4 (Hyperparameter Sensitivity): How do the hyperparameters affect the performance in each experimental datasets? RQ5: How do the imbalance clusters impact the clustering results? Besides, we examine the embedding expressiveness of ASIL, and include a visualization of the clusters learnt from ASIL. Additional results and further details are provided in Appendix VII. Codes and data are publicly available at https://github.com/RiemannGraph/DSE_clustering.